Title

Article text.

-- %USERSIG{BarryRothberg - 2018-05-03}%

Comments

 

UT 01:20 - setup. Dark taken on both LUCIs: LUCI1 0219 LUCI2 0001

UT 01:49 - 5 darks completed. LUCI2 had 5 darks at DIT=150sec (0002-0006) and LUCI1 had 5 darks at DIT=120sec (mistake). Then sent LUCI1 for 5 DARKS at 150 starting with 0225

UT 02:45 - humidity at 85%

UT 03:08 - we closed the dome, humidity rising to limit rapidly

UT 04:36 - Humidity still above 90%

UT 04:45 - Humidity drops to 89.2%

UT 05:09 - Humidity up to 92.5%

UT 08:01 - Humidity drops to 87.6%

UT 08:55 - starting first telluric

UT 09:17 LUCI1 0229-0231 aqu for telluric, 0232-0236 telluric, LUCI2: 0006-0009 aqu and 10-13 telluric spectroscopy on LUCI2

UT09:23 LUCI1 doing MOS aql. next 3 images (2 field images and through mask) for each.

UT 09:38 MOS Science underway starting with LUCI1 0241 LUCI2 0018

UT 10:57 In the last 5 min the seeing started blowing up (4"!) Probably last 2-3 exposures affected if not more (LUCI2 0039,40,41,42) (LUCI1 0262,0263,0264)

UT 11:01 Last MOS exposure LUCI1 02265 LUCI2 0042

UT 11:37 Attempt to observe backup longslit target. Aquisition failed, 3" seeing, twilight starting.

UT 12:12 Did calibrations, done for the night
Topic revision: r8 - 03 May 2018, EiichiEgami
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback